# textwidth=0.8textheight=0.9

### v1

@@ 5.2 Applications
PIKAN variants have been applied to address specific computational challenges, such as resolving high-dimensional, non-linear PDEs in both forward and inverse problems. A non-exhaustive list of recent applications of PIKANs is provided in 4. For example, in the solid mechanics domain, PIKAN has been used to model fracture mechanics problems characterized by displacement singularities and sharp stress gradients. In this context, Wang et al. [53] investigated the Mode III crack problem, which involves intense singularities near the crack tip, and demonstrated that the PIKAN framework successfully captured the steep stress gradients and singular fields with notable accuracy. The results indicate that PIKAN effectively resolves near-tip singularities while maintaining stable convergence and efficient training. Unlike conventional architectures that often struggle with the spectral bias when learning high-frequency features near singular points, PIKAN’s ability to adaptively learn univariate activation functions allows for a more precise representation of the steep gradients inherent to fracture mechanics. Quantitative evaluations against analytical solutions and finite element results show that PIKAN accurately predicts both displacement and stress fields near the crack front (see Fig. 8 (a)).
@@
KAN-PointNet Combined KANs with PointNet to predict incompressible steady-state fluid flow over irregular domains; Utilized Jacobi polynomials in shared KANs; Applied to incompressible laminar steady-state flow over a cylinder with varying cross-sectional geometries. [226]
@@
PI-KAN-PointNet Proposed physics-informed KAN-PointNet for simultaneous inverse solutions over multiple geometries; Employed Jacobi polynomials; Applied to incompressible steady-state fluid flow predicting velocity, pressure, and temperature across 135 domains. [99]
@@
ChebPIKAN Combined KANs with Chebyshev polynomial basis functions and physics-informed loss to solve fluid dynamics PDEs efficiently; Applied to Allen–Cahn, nonlinear Burgers’, 2D Helmholtz, 2D Kovasznay flow, and 2D Navier–Stokes equations. [104]
@@
Legendre-KAN Proposed a neural network framework using Legendre polynomial basis; Applied to high-dimensional and singular fully nonlinear Monge–Ampére equations with Dirichlet boundary conditions and the transport problem (Brenier Theorem). [227]
@@
tanh-cPIKAN Introduced a Chebyshev-based PIKAN variant; Applied to inverse problems in pharmacology systems, including pharmacokinetics (single-dose compartmental model) and pharmacodynamics (chemotherapy drug-response modeling). [102]
@@
Anant-Net Proposed a KAN-based surrogate framework tailored for solving high-dimensional PDEs on hypercubic domains; Employed spline based basis functions; Applied to Poisson, Sine-Gordon, and Allen–Cahn equations up to 300 dimensions. [228]
@@
AIVT Introduced artificial intelligence velocimetry-thermometry for reconstructing temperature and velocity fields; Used Chebyshev KANs; Applied to thermal turbulence in Rayleigh–Bénard convection using experimental 3D velocity data. [229]
@@
ICKANs Proposed monotonic input-convex KANs for learning polyconvex constitutive models in hyperelasticity; Used uniform B-splines; Applied to Neo-Hookean, Isihara, Haines-Wilson, Gent-Thomas, Arruda-Boyce and Ogden models. [230]
@@
Figure 8: (a) The mode III crack problem is defined over a square domain of $[-1,1]^{2}$, with the angular coordinate $\theta$ ranging from $[-\pi,\pi]$. The analytical displacement solution is given by $u={r^{1/2}}\sin(\theta/2)$. This panel is adopted from [53]. (b) The model consists of a square plate with a 20 mm edge length and a centrally located circular hole of 5 mm radius, with predicted displacement solutions shown for PIKAN versions. This panel is adopted from [53]. (c) Description of the Neo-Hookean hyperelastic cantilever beam problem with dimensions of: height $H=1.0$, length $L=4.0$, $E=1000$, and $\nu=0.3$. The left end $x=0.0$ is fixed, and a uniformly distributed downward load $\bar{t}=5$ is applied at the right end. Distributions of absolute errors for the Neo-Hookean hyperelastic cantilever beam obtained from the DEM using the KAN approach. This panel is adopted from [53]. (d) A square plate with two asymmetrically positioned elliptical holes is subjected to uniaxial tension under displacement control. The resulting deformed configuration, with displacement magnitudes represented by color contours, is shown for both the ground-truth data and the ICKAN-based simulation results. This panel is adopted from [230].
@@
In another benchmark by Wang et al. [53], a plate with a central hole was subjected to uniaxial tension to assess the model’s ability to resolve stress concentrations, a crucial test for evaluating the fidelity of surrogate models in capturing localized field amplification. PIKAN demonstrated superior performance in reconstructing both displacement and stress fields in the vicinity of the hole, where stress gradients become highly localized due to geometric discontinuity. PIKAN achieved an accurate approximation of steep solution features with efficient convergence, requiring fewer training epochs. Fig. 8(b) illustrates the initial configuration along with the corresponding solutions obtained using the FEM and PIKAN frameworks. This efficiency stems from the KAN structure’s ability to represent sharp variations in stress distribution without excessive network depth or tuning.
@@
Beyond linear elasticity, the method was extended to simulate nonlinear hyperelastic responses using a Neo-Hookean material model by Wang et al. [53]. Under large deformations and boundary-driven displacement conditions, PIKAN approximated the energy functional and displacement field with high fidelity, yielding results consistent with reference finite element solutions. The variational formulation adopted in this case allowed PIKAN to learn the equilibrium configuration by minimizing the stored strain energy, a strategy that proved effective in capturing nonlinear stress–strain behavior without explicit supervision of the stress field. One notable advantage observed was PIKAN’s robustness in handling geometric and material nonlinearity without the need for architectural adjustments or retraining instability, which are common issues in traditional PINNs under large-deformation regimes. Moreover, the ability of the network to preserve physical constraints such as incompressibility and boundary condition adherence was markedly improved, owing to the localized adaptability of KAN-based activations. Fig. 8(c) presents the Neo-Hookean hyperelastic cantilever beam configuration along with the corresponding absolute error contours obtained from the DEM using the KAN framework.
@@
To show the potential of PIKANs in solid mechanics, Thakolkaran et al. [230] proposed a monotonic Input-Convex KAN (ICKAN) architecture tailored to identify hyperelastic functionals that adhere to polyconvexity constraints. Unlike traditional supervised training on stress–strain data, their model was trained in an unsupervised manner using full-field strain data and global reaction forces, mimicking digital image correlation (DIC) experiments. The ICKAN model enforced convexity with respect to input invariants and monotonicity in the output through structured B-spline activation layers, ensuring thermodynamic consistency and stable optimization. Results demonstrated that the learned energy models accurately captured the underlying physical behavior and generalized well beyond the training distribution, including to unseen geometries and deformation modes. Fig. 8(d) shows finite element solutions for the Isihara and Arruda-Boyce ground-truth models, which closely match the predictions produced by the ICKAN-based models.
@@
In addition to solid mechanics, PIKAN-based frameworks have also been extended to model complex fluid flow phenomena. In this context, Kashefi et al. [99] proposed the PI-KAN-PointNet framework for reconstructing steady-state velocity, pressure, and temperature fields in natural convection problems across 135 two-dimensional domains with internal non-circular cylinders. The network was tasked with inferring flow fields from sparse sensor data and partial boundary conditions, particularly estimating the unknown temperature distribution along the surface of the inner cylinder. Results showed that PI-KAN-PointNet achieved high prediction accuracy, with relative errors below 3% for pressure and temperature fields, and under 11% for velocity components, even in the presence of sharp gradients induced by geometric discontinuities. Fig. 9(a) shows that the physics-informed KAN PointNet accurately reconstructs the velocity and pressure fields across complex geometries, achieving close agreement with the reference solution. The model’s ability to adaptively learn activation functions via Jacobi polynomial expansions proved especially effective near irregular boundaries. Visual comparisons with finite element ground truth confirmed the framework’s strength in capturing localized flow features, especially around re-entrant corners and high-curvature regions of the internal obstacles. Furthermore, the hybrid architecture using MLP encoders and KAN decoders enhanced convergence stability and reduced training epochs, showcasing the benefit of combining geometric feature extraction with physics-informed functional adaptability.
@@
Figure 9: (a) Comparison between the reference solution and the predictions generated by the physics-informed KAN PointNet for the velocity and pressure fields. The model employs Jacobi polynomials of degree 2, with parameters $\alpha=\beta=-0.5$, and uses a sample density of $n_{s}=0.5$. This panel is adopted from [99]. (b) Absolute point-wise testing error evaluated on randomly selected 3D subspaces extracted from the 21D, 99D, and 300D Poisson equation benchmarks. The notation “ActiveDims: {p, q, r}” refers to the dimensions in the $d$-dimensional space where the gradient is non-zero, while $\text{D}\setminus\{p,q,r\}$ indicates the remaining inactive dimensions. Here, $\text{D}=\{1,2,\dots,d\}$ represents the complete set of dimensions. This panel is adopted from [228]. (c) Snapshots illustrate the evolution of the flow over 10 free-fall time units. This panel is adopted from [229]. (d) Diffusion equation: Histogram of predicted values for $D$ obtained using Hamiltonian Monte Carlo (HMC) and Dropout Tikhonov Ensemble Kalman Inversion (DTEKI) with Chebyshev-based KANs. This panel is adopted from [223].
@@
Another notable application of PIKAN-based architectures in fluid modeling is Anant-Net, introduced by Menon & Jagtap [228], which integrates MLPs with KANs to solve high-dimensional nonlinear PDEs efficiently. As a benchmark task, the authors applied Anant-Net to the high-dimensional Poisson equation, a canonical elliptic PDE widely used to evaluate the scalability and numerical stability of surrogate models. The model was tested on domains up to 100 dimensions with manufactured solutions, allowing for quantitative error evaluation. Fig. 9(b) presents the absolute point-wise testing error computed over randomly selected 3D subspaces from the 21D, 99D, and 300D Poisson equation benchmarks. The results demonstrate that Anant-Net maintains low prediction error in the active dimensions, even as the ambient dimensionality increases, showing the scalability and robustness of the proposed architecture in high-dimensional settings.
@@
In another benchmark, Toscano et al. [229] proposed an artificial intelligence velocimetry-thermometry (AIVT) framework based on PIKANs, designed to reconstruct full 3D flow fields in turbulent Rayleigh–Bénard convection from sparse experimental measurements. The AIVT model was trained using sparse velocity measurements obtained via Lagrangian particle tracking, combined with partial temperature data from particle image thermometry. By incorporating the governing equations of momentum and energy, as well as boundary constraints, into the loss function, AIVT inferred smooth, instantaneous flow fields consistent with physical laws. Fig. 9(c) illustrates that the framework accurately reconstructs temperature fluctuation, velocity magnitude, viscosity dissipation, and temperature dissipation rate across space and time, capturing key features such as plume detachment, vorticity intensification, and dissipation zones. Compared to raw measurements, the reconstructed fields exhibited improved spatial coherence and better gradient resolution, particularly in regions lacking direct measurements. The use of KANs enabled localized functional adaptation, mitigating the spectral bias and enhancing the model’s ability to resolve fine-scale turbulent structures.
@@
Expanding the scope of fluid dynamics applications, Gao & Karniadakis [223] introduced a scalable gradient-free PIKAN framework tailored to infer spatially varying diffusion coefficients in elliptic PDEs under sparse and noisy observations. The authors introduced a scaled variant of the method, referred to as SDTEKI, which improves numerical stability and convergence, particularly under data scarcity and high-noise conditions. As a representative benchmark, the model was applied to a two-dimensional steady-state diffusion problem, where sparse and noisy pressure data served as training observations for inferring the underlying diffusion field. The results, drawn from the 1D diffusion equation benchmark, indicate that while all methods successfully recover the true parameter range, SDTEKI yields a more concentrated and stable posterior distribution compared to Hamiltonian Monte Carlo (HMC) and DTEKI, demonstrating improved robustness and inference efficiency under sparse and noisy data conditions. Fig. 9(d) presents histograms of the posterior estimates for the scalar diffusion parameter $D$ obtained using HMC, standard DTEKI, and SDTEKI.
@@ 5.3.1 Accuracy Comparison
Table 5: Comparison of various physics-informed models based on basis functions, optimization strategies, model complexity (number of trainable parameters $|\bm{\theta}|$), and relative $\mathcal{L}^{2}$ error in solving the Allen–Cahn equation. Results are reported across multiple studies with different spatial-temporal domains: $x\in[-1,1],\ t\in[0,1]$ for [66, 117], and $x\in[-6,6],\ t\in[0,1]$ for [68]
@@
PIKAN B-spline Adam 672167216721 $5.83\times 10^{-1}$ [66]
PIKAN Chebyshev Adam 672067206720 $5.15\times 10^{-2}$
PIKAN+RBA Chebyshev Adam 672067206720 $5.65\times 10^{-2}$
PINN+RBA tanh Adam 500495004950049 $1.51\times 10^{-2}$
Scaled-PIKAN Chebyshev Adam 756075607560 $5.90\times 10^{-2}$ [68]
Scaled-PINN tanh Adam 785078507850 $5.10\times 10^{-1}$
PIKAN Chebyshev Adam 756075607560 $6.00\times 10^{-1}$
PINN tanh Adam 785078507850 $5.10\times 10^{-1}$
PIKAN Ku̇rková (KKAN) Adam 195721957219572 $4.41\times 10^{-3}$ [117]
PIKAN+ssRBA Ku̇rková (KKAN) Adam 118491118491118491 $3.07\times 10^{-5}$
PIKAN Chebyshev Adam 203612036120361 $3.55\times 10^{-3}$
PIKAN+ssRBA Chebyshev Adam 104136104136104136 $2.91\times 10^{-4}$
PINN tanh Adam 213182131821318 $1.36\times 10^{-2}$
PINN+ssRBA tanh Adam 915219152191521 $3.52\times 10^{-5}$
PIKAN B-spline Adam + BFGS 554055405540 $1.17\times 10^{-2}$ [107]
PIKAN B-spline Adam + SSBroyden 554055405540 $4.52\times 10^{-4}$
PIKAN Chebyshev Adam + BFGS 136813681368 $1.23\times 10^{-3}$
PIKAN Chebyshev Adam + SSBroyden 136813681368 $9.01\times 10^{-6}$
PINN tanh Adam + BFGS 201920192019 $7.59\times 10^{-4}$
PINN tanh Adam + SSBroyden 201920192019 $1.15\times 10^{-6}$
PINN tanh ssBroyden 201920192019 $1.28\times 10^{-6}$
@@
Evaluating the accuracy of physics-informed models is essential for understanding their ability to approximate complex solutions to PDEs. In the case of PIKAN, a key advantage lies in its use of basis functions, which can offer higher accuracy with fewer trainable parameters compared to vanilla PINNs. Table 5 presents a non-exhaustive overview of studies comparing PIKAN and PINN across various configurations, including different basis types, optimization strategies, and problem settings for solving the Allen-Cahn equation. as an example problem. These comparisons not only highlight the superior accuracy of PIKANs in many cases but also demonstrate how architectural and algorithmic choices influence its effectiveness.
@@
For instance, the choice of optimization strategy plays a critical role in determining the accuracy of physics-informed models. In the work of Kiyani et al. [107], the performance of different optimizers was systematically evaluated for both PIKAN and PINN architectures. The results show that combining the Adam optimizer with SSBroyden (Self-Scaled Broyden) significantly improves the accuracy of PIKAN. This improvement is primarily due to the self-scaling property of SSBroyden, which enables the optimizer to adapt to the local geometry of the loss landscape, thereby enhancing stability and robustness, especially for ill-conditioned problems such as nonlinear PDEs. Specifically, as shown in Table 5, Chebyshev-based PIKAN with hybrid Adam/SSBroyden achieved a 99.3% error reduction compared to the same architecture optimized with hybrid Adam/BFGS. However, achieving this level of accuracy in PINNs, on the order of $10^{-6}$ relative error, required an even greater number of trainable parameters than in Chebyshev-based PIKAN. These findings confirm that PIKAN architectures are more capable of achieving high-accuracy solutions with fewer parameters, particularly when combined with suitable basis functions such as Chebyshev and advanced optimizers like SSBroyden, making them a more practical and efficient choice for solving complex PDEs, such as the Allen–Cahn equation.
@@
Figure 10: Comparison of predicted solutions for the Allen–Cahn equation with diffusion coefficient $D=1\times 10^{-4}$. (a) Solution over the spatio-temporal domain [6,6]×[0,1][-6,6$\times$[0,1][ - 6 , 6 ] × [ 0 , 1 ]. The model was trained using 50,000 residual points, with architectures containing 7,560 CKAN parameters and an MLP with 7,850 parameters. This panel is adapted from [68]. (b) Solution over the spatio-temporal domain [1,1]×[0,1][-1,1$\times$[0,1][ - 1 , 1 ] × [ 0 , 1 ]. The network used 50000 residual points, with 6720 KAN parameters and an MLP with 50049 parameters. This panel is adopted from [66].

### v2

@@ 5.2 Applications
PIKAN variants have been applied to address specific computational challenges, such as resolving high-dimensional, non-linear PDEs in both forward and inverse problems. A non-exhaustive list of recent applications of PIKANs is provided in Table 4. For example, in the solid mechanics domain, PIKAN has been used to model fracture mechanics problems characterized by displacement singularities and sharp stress gradients. In this context, Wang et al. wang2024kolmogorov investigated the Mode III crack problem, which involves intense singularities near the crack tip, and demonstrated that the PIKAN framework successfully captured the steep stress gradients and singular fields with notable accuracy. The results indicate that PIKAN effectively resolves near-tip singularities while maintaining stable convergence and efficient training. Unlike conventional architectures that often struggle with the spectral bias when learning high-frequency features near singular points, PIKAN’s ability to adaptively learn univariate activation functions allows for a more precise representation of the steep gradients inherent to fracture mechanics. Quantitative evaluations against analytical solutions and finite element results show that PIKAN accurately predicts both displacement and stress fields near the crack front (see Fig. 8 (a)).
@@
KAN-PointNet Combined KANs with PointNet to predict incompressible steady-state fluid flow over irregular domains; Utilized Jacobi polynomials in shared KANs; Applied to incompressible laminar steady-state flow over a cylinder with varying cross-sectional geometries. KASHEFI2025117888
@@
PI-KAN-PointNet Proposed physics-informed KAN-PointNet for simultaneous inverse solutions over multiple geometries; Employed Jacobi polynomials; Applied to incompressible steady-state fluid flow predicting velocity, pressure, and temperature across 135 domains. kashefi2025physics
@@
ChebPIKAN Combined KANs with Chebyshev polynomial basis functions and physics-informed loss to solve fluid dynamics PDEs efficiently; Applied to Allen–Cahn, nonlinear Burgers’, 2D Helmholtz, 2D Kovasznay flow, and 2D Navier–Stokes equations. guo2024physics
@@
Legendre-KAN Proposed a neural network framework using Legendre polynomial basis; Applied to high-dimensional and singular fully nonlinear Monge–Ampére equations with Dirichlet boundary conditions and the transport problem (Brenier Theorem). hu2025s
@@
tanh-cPIKAN Introduced a Chebyshev-based PIKAN variant; Applied to inverse problems in pharmacology systems, including pharmacokinetics (single-dose compartmental model) and pharmacodynamics (chemotherapy drug-response modeling). daryakenari2025representation
@@
Anant-Net Proposed a KAN-based surrogate framework tailored for solving high-dimensional PDEs on hypercubic domains; Employed spline based basis functions; Applied to Poisson, Sine-Gordon, and Allen–Cahn equations up to 300 dimensions. menon2025
@@
AIVT Introduced artificial intelligence velocimetry-thermometry for reconstructing temperature and velocity fields; Used Chebyshev KANs; Applied to thermal turbulence in Rayleigh–Bénard convection using experimental 3D velocity data. toscano2025aivt
@@
ICKANs Proposed monotonic input-convex KANs for learning polyconvex constitutive models in hyperelasticity; Used uniform B-splines; Applied to Neo-Hookean, Isihara, Haines-Wilson, Gent-Thomas, Arruda-Boyce and Ogden models. thakolkaran2025can
@@
Figure 8: (a) The mode III crack problem is defined over a square domain of $1,1]^{2}$, with the angular coordinate $\theta$ ranging from $\pi,\pi]$. The analytical displacement solution is given by $u={r^{1/2}}\sin(\theta/2)$. This panel is adopted from wang2024kolmogorov . (b) The model consists of a square plate with a 20 mm edge length and a centrally located circular hole of 5 mm radius, with predicted displacement solutions shown for PIKAN versions. This panel is adopted from wang2024kolmogorov . (c) Description of the Neo-Hookean hyperelastic cantilever beam problem with dimensions of: height $H=1.0$, length $L=4.0$, $E=1000$, and $\nu=0.3$. The left end $x=0.0$ is fixed, and a uniformly distributed downward load $\bar{t}=5$ is applied at the right end. Distributions of absolute errors for the Neo-Hookean hyperelastic cantilever beam obtained from the DEM using the KAN approach. This panel is adopted from wang2024kolmogorov . (d) A square plate with two asymmetrically positioned elliptical holes is subjected to uniaxial tension under displacement control. The resulting deformed configuration, with displacement magnitudes represented by color contours, is shown for both the ground-truth data and the ICKAN-based simulation results. This panel is adopted from thakolkaran2025can .
@@
In another benchmark by Wang et al. wang2024kolmogorov , a plate with a central hole was subjected to uniaxial tension to assess the model’s ability to resolve stress concentrations, a crucial test for evaluating the fidelity of surrogate models in capturing localized field amplification. PIKAN demonstrated superior performance in reconstructing both displacement and stress fields in the vicinity of the hole, where stress gradients become highly localized due to geometric discontinuity. PIKAN achieved an accurate approximation of steep solution features with efficient convergence, requiring fewer training epochs. Fig. 8(b) illustrates the initial configuration along with the corresponding solutions obtained using the FEM and PIKAN frameworks. This efficiency stems from the KAN structure’s ability to represent sharp variations in stress distribution without excessive network depth or tuning.
@@
Beyond linear elasticity, the method was extended to simulate nonlinear hyperelastic responses using a Neo-Hookean material model by Wang et al. wang2024kolmogorov . Under large deformations and boundary-driven displacement conditions, PIKAN approximated the energy functional and displacement field with high fidelity, yielding results consistent with reference finite element solutions. The variational formulation adopted in this case allowed PIKAN to learn the equilibrium configuration by minimizing the stored strain energy, a strategy that proved effective in capturing nonlinear stress–strain behavior without explicit supervision of the stress field. One notable advantage observed was PIKAN’s robustness in handling geometric and material nonlinearity without the need for architectural adjustments or retraining instability, which are common issues in traditional PINNs under large-deformation regimes. Moreover, the ability of the network to preserve physical constraints such as incompressibility and boundary condition adherence was markedly improved, owing to the localized adaptability of KAN-based activations. Fig. 8(c) presents the Neo-Hookean hyperelastic cantilever beam configuration along with the corresponding absolute error contours obtained from the DEM using the KAN framework.
@@
To show the potential of PIKANs in solid mechanics, Thakolkaran et al. thakolkaran2025can proposed a monotonic Input-Convex KAN (ICKAN) architecture tailored to identify hyperelastic functionals that adhere to polyconvexity constraints. Unlike traditional supervised training on stress–strain data, their model was trained in an unsupervised manner using full-field strain data and global reaction forces, mimicking digital image correlation (DIC) experiments. The ICKAN model enforced convexity with respect to input invariants and monotonicity in the output through structured B-spline activation layers, ensuring thermodynamic consistency and stable optimization. Results demonstrated that the learned energy models accurately captured the underlying physical behavior and generalized well beyond the training distribution, including to unseen geometries and deformation modes. Fig. 8(d) shows finite element solutions for the Isihara and Arruda-Boyce ground-truth models, which closely match the predictions produced by the ICKAN-based models.
@@
In addition to solid mechanics, PIKAN-based frameworks have also been extended to model complex fluid flow phenomena. In this context, Kashefi et al. kashefi2025physics proposed the PI-KAN-PointNet framework for reconstructing steady-state velocity, pressure, and temperature fields in natural convection problems across 135 two-dimensional domains with internal non-circular cylinders. The network was tasked with inferring flow fields from sparse sensor data and partial boundary conditions, particularly estimating the unknown temperature distribution along the surface of the inner cylinder. Results showed that PI-KAN-PointNet achieved high prediction accuracy, with relative errors below 3% for pressure and temperature fields, and under 11% for velocity components, even in the presence of sharp gradients induced by geometric discontinuities. Fig. 9(a) shows that the physics-informed KAN PointNet accurately reconstructs the velocity and pressure fields across complex geometries, achieving close agreement with the reference solution. The model’s ability to adaptively learn activation functions via Jacobi polynomial expansions proved especially effective near irregular boundaries. Visual comparisons with finite element ground truth confirmed the framework’s strength in capturing localized flow features, especially around re-entrant corners and high-curvature regions of the internal obstacles. Furthermore, the hybrid architecture using MLP encoders and KAN decoders enhanced convergence stability and reduced training epochs, showcasing the benefit of combining geometric feature extraction with physics-informed functional adaptability.
@@
Figure 9: (a) Comparison between the reference solution and the predictions generated by the physics-informed KAN PointNet for the velocity and pressure fields. The model employs Jacobi polynomials of degree 2, with parameters $\alpha=\beta=-0.5$, and uses a sample density of $n_{s}=0.5$. This panel is adopted from kashefi2025physics . (b) Absolute point-wise testing error evaluated on randomly selected 3D subspaces extracted from the 21D, 99D, and 300D Poisson equation benchmarks. The notation “ActiveDims: {p, q, r}” refers to the dimensions in the $d$-dimensional space where the gradient is non-zero, while $\text{D}\setminus\{p,q,r\}$ indicates the remaining inactive dimensions. Here, $\text{D}=\{1,2,\dots,d\}$ represents the complete set of dimensions. This panel is adopted from menon2025 . (c) Snapshots illustrate the evolution of the flow over 10 free-fall time units. This panel is adopted from toscano2025aivt . (d) Diffusion equation: Histogram of predicted values for $D$ obtained using Hamiltonian Monte Carlo (HMC) and Dropout Tikhonov Ensemble Kalman Inversion (DTEKI) with Chebyshev-based KANs. This panel is adopted from gao2025scalable .
@@
Another notable application of PIKAN-based architectures in fluid modeling is Anant-Net, introduced by Menon & Jagtap menon2025 , which integrates MLPs with KANs to solve high-dimensional nonlinear PDEs efficiently. As a benchmark task, the authors applied Anant-Net to the high-dimensional Poisson equation, a canonical elliptic PDE widely used to evaluate the scalability and numerical stability of surrogate models. The model was tested on domains up to 100 dimensions with manufactured solutions, allowing for quantitative error evaluation. Fig. 9(b) presents the absolute point-wise testing error computed over randomly selected 3D subspaces from the 21D, 99D, and 300D Poisson equation benchmarks. The results demonstrate that Anant-Net maintains low prediction error in the active dimensions, even as the ambient dimensionality increases, showing the scalability and robustness of the proposed architecture in high-dimensional settings.
@@
In another benchmark, Toscano et al. toscano2025aivt proposed an artificial intelligence velocimetry-thermometry (AIVT) framework based on PIKANs, designed to reconstruct full 3D flow fields in turbulent Rayleigh–Bénard convection from sparse experimental measurements. The AIVT model was trained using sparse velocity measurements obtained via Lagrangian particle tracking, combined with partial temperature data from particle image thermometry. By incorporating the governing equations of momentum and energy, as well as boundary constraints, into the loss function, AIVT inferred smooth, instantaneous flow fields consistent with physical laws. Fig. 9(c) illustrates that the framework accurately reconstructs temperature fluctuation, velocity magnitude, viscosity dissipation, and temperature dissipation rate across space and time, capturing key features such as plume detachment, vorticity intensification, and dissipation zones. Compared to raw measurements, the reconstructed fields exhibited improved spatial coherence and better gradient resolution, particularly in regions lacking direct measurements. The use of KANs enabled localized functional adaptation, mitigating the spectral bias and enhancing the model’s ability to resolve fine-scale turbulent structures.
@@
Expanding the scope of fluid dynamics applications, Gao & Karniadakis gao2025scalable introduced a scalable gradient-free PIKAN framework tailored to infer spatially varying diffusion coefficients in elliptic PDEs under sparse and noisy observations. The authors introduced a scaled variant of the method, referred to as SDTEKI, which improves numerical stability and convergence, particularly under data scarcity and high-noise conditions. As a representative benchmark, the model was applied to a two-dimensional steady-state diffusion problem, where sparse and noisy pressure data served as training observations for inferring the underlying diffusion field. The results, drawn from the 1D diffusion equation benchmark, indicate that while all methods successfully recover the true parameter range, SDTEKI yields a more concentrated and stable posterior distribution compared to Hamiltonian Monte Carlo (HMC) and DTEKI, demonstrating improved robustness and inference efficiency under sparse and noisy data conditions. Fig. 9(d) presents histograms of the posterior estimates for the scalar diffusion parameter $D$ obtained using HMC, standard DTEKI, and SDTEKI.
@@ 5.3.1 Accuracy Comparison
Table 5: Comparison of various physics-informed models based on basis functions, optimization strategies, model complexity (number of trainable parameters $|\bm{\theta}|$), and relative $\mathcal{L}^{2}$ error in solving the Allen–Cahn equation. Results are reported across multiple studies with different spatial-temporal domains: $x\in[-1,1],\ t\in[0,1]$ for shukla2024comprehensive , toscano2025kkans , and $x\in[-6,6],\ t\in[0,1]$ for mostajeran2025scaled
@@
PIKAN B-spline Adam $6721$ $5.83\times 10^{-1}$ shukla2024comprehensive
PIKAN Chebyshev Adam $6720$ $5.15\times 10^{-2}$
PIKAN+RBA Chebyshev Adam $6720$ $5.65\times 10^{-2}$
PINN+RBA tanh Adam $50049$ $1.51\times 10^{-2}$
Scaled-PIKAN Chebyshev Adam $7560$ $5.90\times 10^{-2}$ mostajeran2025scaled
Scaled-PINN tanh Adam $7850$ $5.10\times 10^{-1}$
PIKAN Chebyshev Adam $7560$ $6.00\times 10^{-1}$
PINN tanh Adam $7850$ $5.10\times 10^{-1}$
PIKAN Ku̇rková (KKAN) Adam $19572$ $4.41\times 10^{-3}$ toscano2025kkans
PIKAN+ssRBA Ku̇rková (KKAN) Adam $118491$ $3.07\times 10^{-5}$
PIKAN Chebyshev Adam $20361$ $3.55\times 10^{-3}$
PIKAN+ssRBA Chebyshev Adam $104136$ $2.91\times 10^{-4}$
PINN tanh Adam $21318$ $1.36\times 10^{-2}$
PINN+ssRBA tanh Adam $91521$ $3.52\times 10^{-5}$
PIKAN B-spline Adam + BFGS $5540$ $1.17\times 10^{-2}$ kiyani2025optimizer
PIKAN B-spline Adam + SSBroyden $5540$ $4.52\times 10^{-4}$
PIKAN Chebyshev Adam + BFGS $1368$ $1.23\times 10^{-3}$
PIKAN Chebyshev Adam + SSBroyden $1368$ $9.01\times 10^{-6}$
PINN tanh Adam + BFGS $2019$ $7.59\times 10^{-4}$
PINN tanh Adam + SSBroyden $2019$ $1.15\times 10^{-6}$
PINN tanh ssBroyden $2019$ $1.28\times 10^{-6}$
@@
Evaluating the accuracy of physics-informed models is essential for understanding their ability to approximate complex solutions to PDEs. In the case of PIKAN, a key advantage lies in its use of basis functions, which can offer higher accuracy with fewer trainable parameters compared to vanilla PINNs. Table 5 presents a non-exhaustive overview of studies comparing PIKAN and PINN across various configurations, including different basis types, optimization strategies, and problem settings for solving the Allen-Cahn equation as an example problem. The effect of hyperparameter selection, such as polynomial degree, network configuration, and the number of training samples, on the accuracy of the PIKAN model is also discussed. These comparisons not only highlight the superior accuracy of PIKANs in many cases but also demonstrate how architectural and algorithmic choices influence its effectiveness.
@@
For instance, the choice of optimization strategy plays a critical role in determining the accuracy of physics-informed models. In the work of Kiyani et al. kiyani2025optimizer , the performance of different optimizers was systematically evaluated for both PIKAN and PINN architectures. The results show that combining the Adam optimizer with SSBroyden (Self-Scaled Broyden) significantly improves the accuracy of PIKAN. This improvement is primarily due to the self-scaling property of SSBroyden, which enables the optimizer to adapt to the local geometry of the loss landscape, thereby enhancing stability and robustness, especially for ill-conditioned problems such as nonlinear PDEs. Specifically, as shown in Table 5, Chebyshev-based PIKAN with hybrid Adam/SSBroyden achieved a 99.3% error reduction compared to the same architecture optimized with hybrid Adam/BFGS. However, achieving this level of accuracy in PINNs, on the order of $10^{-6}$ relative error, required an even greater number of trainable parameters than in Chebyshev-based PIKAN. These findings confirm that PIKAN architectures are more capable of achieving high-accuracy solutions with fewer parameters, particularly when combined with suitable basis functions such as Chebyshev and advanced optimizers like SSBroyden, making them a more practical and efficient choice for solving complex PDEs, such as the Allen–Cahn equation.
@@
Figure 10: Comparison of predicted solutions for the Allen–Cahn equation with diffusion coefficient $D=1\times 10^{-4}$. (a) Solution over the spatio-temporal domain $6,6]\times[0,1]$. The model was trained using 50,000 residual points, with architectures containing 7,560 CKAN parameters and an MLP with 7,850 parameters. This panel is adapted from mostajeran2025scaled . (b) Solution over the spatio-temporal domain $1,1]\times[0,1]$. The network used 50000 residual points, with 6720 KAN parameters and an MLP with 50049 parameters. This panel is adopted from shukla2024comprehensive .